OpenStudy (anonymous):

What is the slope of the line that is perpendicular to y = -5x + 13 ?

OpenStudy (anonymous):

well use y=mx+b, and remember that parallel lines have the same slope. so what is the slope? :)

OpenStudy (anonymous):

@860bu2y are you still here?

OpenStudy (anonymous):

is it 1/5?

OpenStudy (anonymous):

two lines are perpendicular if the product of their slopes =-1

OpenStudy (anonymous):

oh darn misread the question XD

OpenStudy (anonymous):

using y=mx+c , m=-5

OpenStudy (anonymous):

let m2 be the slope of our line...then m*m2 =-1 -5*m2=-1 m2 =1/5

OpenStudy (anonymous):

so the answer is 1/5